3. Cam, Inc. currently sells widgets for $80 per unit. The variable cost is $30 per unit and total fixed costs equal $240,000 per year. Sales are currently 20,000 units annually Required Calculate the bontribution margin ratio (round your answer to xxx%). b. Calculate break-even in units. c. Calculate the break-even in sales dollars d. Calculate the current operating income assuming no income taxes. The company is considering a 20% drop in selling price that it believes will raise units sold by 20%. all costs stay the same, what is the impact on income if this chsange is made? Assuming
Answer #1

A:- contribution margin ratio

Sales - variable cost / sales x 100 = 80 - 30 / 80 x 100 = 62.5 %

B:- break even sales in units

Fixed cost / sales - variable cost

= 240,000/ 80-30

= 4800 units

C :- break even sales in doller

= break even in units x sales per unit

= 4800 x 80 = 384,000

D:- Operating income

= sales - variable cost - fixed cost

= 20,000 x 80 - 20,000 x 3 - 240,000

= 16,00,000 - 6,00,000 - 240,000

= 760,000

E:- Current Scenario:

SP $80

VC 30

CM $50 × 20,000 units = $1,000,000 total contribution margin

Proposed Change:

SP $80 ×80% = $64

VC 30

CM $34 ×(20,000 ×1.2) = $816,000 total contribution margin Total CM (and income) drops by $184,000 ($1,000,000 – $816,000).